OpenStudy (johnnewman):

If f(x) = 0.5(1 – x), what is the value of f(–1)? A. 2.5 B. 1 C. 0.5 D. 0

OpenStudy (plainntall):

It is asking you to evaluate the function at x=-1 so just substitute and do the math 0.5(1-(-1))=0.5(1+1)=0.5(2)=?

OpenStudy (johnnewman):

B.

OpenStudy (plainntall):

Yes :D In problems like this just substitute the value and do the math.